OpenStudy (anonymous):

Express f(x) in the form f(x) = A Sin (bx+c) for the function: f(x) = sin (x) + sqrt3 cos (x)

OpenStudy (anonymous):

\[disambiguation: f(x) = \sin(x) + \sqrt{3}\cos (x) \]

OpenStudy (anonymous):

alright, i really have no way to explain it, it was a guess and check on my part, but I got: \[f(x) = 2\sin(x+\frac{\pi}{3})\] expanding this using sum of two angle formulas gives:\[2\sin(x+\frac{\pi}{3})=2(\sin(x)(\frac{1}{2})+\cos(x)(\frac{\sqrt3}{2}))\] which simplifies back to what you started with.

OpenStudy (anonymous):

\[A \sin(x + \phi) = A \sin x \cos \phi + A \cos x \sin \phi = \sqrt{3} \cos x\] for all x. Hence, \[A \cos \phi = 1\]\[ A \sin \phi = \sqrt{3}.\]We have, then that \[ \tan \phi = \sqrt{3},\]so \[\phi = \pi /3.\]If we square both sides of both equations and then add, we get \[(A \cos \phi) ^{2} + (A \sin \phi)^{2} = 4,\]so A = 2, and the answer is \[2 \sin(x + \pi/3).\]

OpenStudy (anonymous):

\[f(x) = 2 \sin (x+\pi /3)\] is the answer given in the key. I very much appreciate the work! I will work to decipher the steps.

OpenStudy (anonymous):

Right on:^)